OpenStudy (hahawowoh):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part A: Write functions to represent Cory and Roger's collections throughout the years. (4 points) Part B: How many cars does Cory have after 6 years? How many does Roger have after the same number of years? (2 points) Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically. (4 points)

OpenStudy (hahawowoh):

please help

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you are not getting an answer because there is no information with which to answer it

OpenStudy (hahawowoh):

there I fixed it

OpenStudy (misty1212):

to increase a number by \(20\%\) multiply it by \(100\%+20\%=120\%=1.2\) so a good function for cory is \[15\times 1.2^t\]

OpenStudy (hahawowoh):

I have everything just part c

OpenStudy (misty1212):

i guess you are supposed to solve \[t+40=15\times (1.2)^t\]

OpenStudy (misty1212):

it says "approximately" so i would use this http://www.wolframalpha.com/input/?i=t%2B40%3D15%281.2%29%5Et

OpenStudy (misty1212):

or since you were already asked how many each had in six years, and since those numbers are almost equal, say "six"